ROTATIONAL SPECTROSCOPY UNVEILS ELEVEN CONFORMERS OF ADRENALINE

C. CABEZAS, V. CORTIJO, S. MATA, J. C. LÓPEZ, \underline{J. L. ALONSO}, Grupo de Espectroscopía Molecular (GEM), Edificio Quifima, Laboratorios de Espectroscopía y Bioespectroscopía, Parque Científico, Universidad de Valladolid, 47011 Valladolid, Spain.

Recent improvements in our LA-MB-FTMW instrumentation have allowed the characterization of eleven and eight conformers for the neurotransmitters adrenaline and noradrenaline respectively. The observation of this rich conformational behavior is in accordance with the recent observation of seven conformers for dopamine^{a} and in sharp contrast with the conformational reduction proposed for catecholamines.^{b}

aC. Cabezas, I. Peña, J. C. López, J. L. Alonso J. Phys. Chem. Lett. 2013, 4, 486.

bH. Mitsuda, M. Miyazaki, I. B. Nielsen, P. arabal,C. Dedonder, C. Jouvet, S. Ishiuchi, M. Fujii J. Phys. Chem. Lett. 2010, 1,